A venture capitalist, willing to invest $1,000,000, has three investments to choose from. The first investment, a software company, has a 10% chance of returning $5,000,000 profit, a 30% chance of returning $1,000,000 profit, and a 60% chance of losing the million dollars. The second company, a hardware company, has a 20% chance of returning $3,000,000 profit, a 40% chance of returning $1,000,000 profit, and a 40% chance of losing the million dollars. The third company, a biotech firm, has a 10% chance of returning $6,000,000 profit, a 70% of no profit or loss, and a 20% chance of losing the million dollars.

Order the expected values from smallest to largest.

1)Biotech, Hardware, Software

2)Biotech, Software, Hardware

3)Software, Hardware, Biotech

4)Hardware, Software, Biotech

5)Software, Biotech, Hardware

6)Hardware, Biotech, Software
Answer #1

expected value of  software company =(0.1*5+0.3*1+0.6*(-1)) =0.2 million

expected value of   hardware company =(0.2*3+0.4*1+0.4*(-1)) =0.6 million

expected value of  biotech firm =(0.1*6+0.7*0+0.2*(-1)) =0.4 million

expected values from smallest to largest:

5)Software, Biotech, Hardware